OpenStudy (anonymous):

PLEASE HELP HOMEWORK DUE TOMORROW perimeter is 384 yards, length is 4 yards less than triple the width. what is the dimensions.

OpenStudy (crashonce):

ok lets set the width as x yards

OpenStudy (crashonce):

therefore the length will be 3x-4

OpenStudy (crashonce):

solve the rest

OpenStudy (anonymous):

i honestly have no skills in math could i have the answers please

OpenStudy (mathmate):

Let W=Width L=length=3*W-4 Perimeter = 2(L+W)=2(3*W-4 + W) = 384 Solve for W (and then L)

OpenStudy (mathmate):

Sorry @Schoolninja7 , we are not allowed to give answers only. If you as you said have no skills in math, that exactly why you're here, and at school, to acquire the necessary skills for your future projects.

OpenStudy (anonymous):

how do i find the wisth

OpenStudy (mathmate):

See the perimeter line? There is an equation involving only one single unknown, W. You need to solve for W and then find the length L=3W-4.

OpenStudy (anonymous):

im still stuck thanks for the help anyway.

OpenStudy (mathmate):

You'd need to solve for W in 2(3*W-4 + W) = 384 First distribute the multiplier "2", 6W-8+2W=384 Then group variables to the left and numbers to the right of the equal sign, then solve for W.

OpenStudy (anonymous):

98

OpenStudy (anonymous):

290 L 98 W

OpenStudy (mathmate):

98 sounds a little big for me. How did you solve it?

OpenStudy (anonymous):

solving multi step equation method

OpenStudy (mathmate):

We have 6W-8+2W=384 so 8W=392 W=49 Can you find L?

OpenStudy (anonymous):

143

OpenStudy (mathmate):

Yes, L=143. (It's always good to say what the number represents). ok, so you're done! L and W are the dimensions!

OpenStudy (anonymous):

thanks a lot

OpenStudy (mathmate):

You're welcome! :)